OpenStudy (anonymous):

Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps.

OpenStudy (math&ing001):

There's infinite forms you can give of x^2. Exemple: \[x ^{3}*x ^{-1}=x ^{3-1}=x ^{2}\]\[\sqrt[3]{x ^{6}}=\sqrt[3]{(x ^{2})^{3}}=x ^{2}\] Check this website and try making the 3rd exemple http://web.mnstate.edu/jamesju/Spr2011/Content/MD127RatExp&Radicals.pdf